OpenStudy (anonymous):

can anyone please help? the revenue from selling x units of a product at a price of p dollars per unit is given by xp. For a particular commodity, the revenue is R=800x-0.25x^2 Factor the expression for the revenue and determine an expression for the price in terms of x.

OpenStudy (anonymous):

help :O

OpenStudy (anonymous):

R=-0.25x^2+800x R=-0.25x(x-32) This will be the factored expression.

OpenStudy (anonymous):

1 sec,

OpenStudy (anonymous):

u sure brah, that looks like an 8 not an 800...

OpenStudy (anonymous):

\[R=5x(40-0.5x)\]

OpenStudy (anonymous):

I think that is better.

OpenStudy (anonymous):

is that the expression for x?

OpenStudy (anonymous):

its says to check the syntax of my response

OpenStudy (anonymous):

Its the expression in terms of x, as your question states, yes.

OpenStudy (anonymous):

According to the problem, it is an 800: R=800x-0.25x^2 40*5 only equals 200. And 5*0.5 equals 2.5.. This is not factoring... R=xp R=-0.25x^2+800x -0.25x(x-32) = xp Do you get it? Now solve for P.

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

So you/I have factored out R=-0.25x^2+800x to be R=-0.25x(x-32). Do you get this step? The problem is asking you to find "the price in terms of x". Your problem tells you that "the revenue from selling x units of a product at a price of p dollars per unit is given by xp". This means R (the revenue) = x*p You also know that R=-0.25x(x-32) (from your factored equation) and R=x*p Since these 2 equations both equal R, that means they must be equal to each other. So, -0.25x(x-32) = xp. You want to solve for P. -0.25x(x-32) = xp Divide both sides by x. -0.25(x-32) = P This is the answer for the problem. You can distribute the -0.25 to the x and -32 if you want.

OpenStudy (anonymous):

help me

OpenStudy (anonymous):

ohhhhhh i see now

OpenStudy (anonymous):

Wait, I think there might be an error.

OpenStudy (anonymous):

-0.25(x-3200)

OpenStudy (anonymous):

i think so :P

OpenStudy (anonymous):

-0.25x(x-3200) is the correct factored form... So replace -0.25x(x-32) with -0.25x(x-3200). Sorry about that.

OpenStudy (anonymous):

-0.25(x-3200) = P

OpenStudy (anonymous):

would i write that as P=(-0.25x)800???

OpenStudy (anonymous):

P = -0.25x+800

OpenStudy (anonymous):

it keeps saying check for the syntax of your responce D:

OpenStudy (anonymous):

What are you doing with it? Is this for online school?

OpenStudy (anonymous):

wouldnt it be -0.25x^2

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

You already canceled out the x when you tried to get P alone.

OpenStudy (anonymous):

so how should i type it in?

OpenStudy (anonymous):

It's asking for 2 things. 1) R = -0.25x(x-3200) 2) P = -0.25x+800

OpenStudy (anonymous):

let me try it

OpenStudy (anonymous):

i give up